# Exponential quantum advantages in learning quantum observables from classical data

### v1

@@ Abstract
Quantum computers are believed to bring computational advantages in simulating quantum many body systems. However, recent works have shown that classical machine learning algorithms are able to predict numerous properties of quantum systems with classical data. Despite various examples of learning tasks with provable quantum advantages being proposed, they all involve cryptographic functions and do not represent any physical scenarios encountered in laboratory settings. In this paper we prove quantum advantages for the physically relevant task of learning quantum observables from classical (measured out) data. We consider two types of observables: first we prove a learning advantage for linear combinations of Pauli strings, then we extend the result for the broader case of unitarily parametrized observables. For each type of observable we delineate the boundaries that separate physically relevant tasks which classical computers can solve using data from quantum measurements, from those where a quantum computer is still necessary for data analysis. Our results shed light on the utility of quantum computers for machine learning problems in the domain of quantum many body physics, thereby suggesting new directions where quantum learning improvements may emerge.
1 Introduction
@@
The very first proposed application of quantum computers can be traced back to Feynman’s idea of simulating quantum physics on a quantum device. Together with factoring, simulation of quantum many body systems stand as the clearest example of dramatic advantages of quantum computers [Llo96]. Machine learning is another much newer area where quantum computers are believed to possibly bring advantages in certain learning problems, and in fact there are provable speed ups achieved by a quantum algorithm [LAT21, SG04, SSHE21, JGM+21] for specific machine learning problems. Relating back to the original Feynman’s idea of simulating quantum physics, machine learning problems in quantum many body physics seem a natural scenario where learning advantages could arise. However, perhaps surprisingly, it was recently shown that access to data exemplifying what the hard-to-compute function does can drastically change the hardness of the computational task questioning the role of quantum computation in a machine learning scenarios [HBM+21, HKT+22b]. If every quantum computation could be replicated classically provided access to data, such results would confine the practical application of quantum computers solely to the data acquisition stage. This is however not the case, as was shown already in the examples considered in [SG04, LAT21]. In such cases the unknown function was cryptographic in nature, and not related to genuine quantum simulation problems. Nonetheless in [GD23] it was elucidated that learning problems with provable speed-ups can be derived from any BQP-complete functions, facilitating connections to more physical scenarios.
The work in [GD23] demonstrated the existence of learning problems with provable advantages based on such reasonings, however it does not provide any natural learning setting of the kind encountered in physics. Another significant open question, raised by the recent progress on classical algorithms with provable guarantees, pertains to establishing clear boundaries between classical and quantum algorithms when dealing with data generated by quantum processes. In this work we provide substantial strides in both directions. We specifically consider learning problems where an experimentalist wants to predict expectation values of an unknown observable from measurements on input quantum states which can either be ground states of local Hamiltonians or time evolved states. Our result is a proof of learning advantages for different types of quantum observables. Firstly, we prove an exponential quantum advantage in learning observables constructed as an unknown linear combination of Pauli strings. Then we extend our quantum speed-up result to the far more general case of unitarily-parametrized observables, linking our results to the extensive body of literature on learning unknown unitaries from query access. In connection to this, we examine the Hamiltonian learning problem where the identification of the target function is demonstrated to be classically easy. We summarize our main results in Table 1.
@@
Linear combination of Pauli strings Time Evolution Problem Yes
@@
Table 1: Main results of this paper
2 Related work
@@
Our main results are built upon the theoretical assumption that there exists a BQP language and an input distribution such that the language cannot be decided using $\mathrm{HeurP/poly}$, for more on this see [GD23]. Building on this, the authors also proved that many learning speed-ups could arise by looking at BQP complete problems. However, in their separation result, the concept class is limited to be of polynomial size, and the quantum algorithm crucially leverages this to find the underlying concept in polynomial time by brute force. In this work we take a step further from their result by considering a continuous concept class and using a quantum algorithm which learns the unknown concept from data through a natural machine learning method.
@@
In recent works by [HKT+22b, LHT+24], significant progresses have been made with classical machine learning algorithms with provable guarantees for learning ground state properties from data, restricting the role of a quantum computer only to the data acquisition phase. In this study, we establish rigorous limitations on the capabilities of classical algorithms when learning quantum observables. A key distinction from their approach is that, in considering ground state problems, we investigate Hamiltonians with a polynomially decaying spectral gap, whereas the algorithms in [HKT+22b, LHT+24] have provable guarantees only for gapped Hamiltonians
@@
In [YIM23] the authors proposed a family of supervised learning tasks which present a learning speed-up with an exponentially sized concept class. Our construction, which was developed independently from [YIM23], stems from a physically-motivated context, resulting in the utilization of different techniques for the quantum learning algorithm given the necessity to handle errors in the training samples. The result of Theorem 3 further differentiates our work, as it deals with unitarily parametrized observables for which the corresponding learning algorithm cannot be straightforwardly simplified to solving a linear system of equations, as it is in the case of [YIM23].
3 The learning problems
@@
We now take in consideration the first scenario where experimentalists can time evolve arbitrary input states under a fixed Hamiltonian for constant time, however they do not have full control over the input state preparation. We model this by allowing the input states to be sampled randomly from an unknown distribution of quantum states, subsequently they undergo a time evolution before being measured by an unknown observable. The experimentalists collect the corresponding measurement outcomes and the goal is to predict expectation values of the unknown observable on new input states. This simple scenario is later significantly generalized to cases where the inputs are known Hamiltonians, and the problem is not time evolution but rather learning measurements on their ground states. In both cases we rigorously model the colloquially described learning tasks using concept classes within the PAC learning framework
Let $H$ be a Hamiltonian and fix a constant time $\tau$, we model the time evolution learning problem by the following concept class
$\mathcal{F}_{\mathrm{evolved}}^{H,O}$ =$=\{f^{\bm{{\bm{\alpha}}}}(\bm{x})\in\mathbb{R}\;\;|\;\mathbf{{\bm{\alpha}}}\in[-1,1]^{m}\}$ (1)
with $\text{with }\;f^{\bm{\alpha}}(\bm{x}):\;\;$ $\;\bm{x}\in\{0,1\}^{n}\rightarrow f^{\bm{\alpha}}(\bm{x})=\text{Tr}[\rho_{H}(\bm{x})O(\mathbf{{\bm{\alpha}}})]\ \;\;\&\;\;O({\bm{\alpha}})=\sum_{i=1}^{m}{\bm{\alpha}}_{i}P_{i}.$ (2)
@@
In the above, $\bm{x}$ specifies the initial state |$\ket{\bm{x}}$, $\rho_{H}(\bm{x})$ the constant time evolved state $\rho_{H}(\bm{x})=U\ket{\bm{x}}\bra{\bm{x}}U^{\dagger}$ with $U=e^{iH\tau}$ and each $P_{i}$ is a $k$-local Pauli string where $m$ scales polynomially with $n$. Notice that although we consider binary inputs $\bm{x}\in\{0,1\}^{n}$, the output of each concept is a real. The goal of the ML learning algorithm is to learn a model $h(\bm{x})$ which approximates the unknown concept $f^{\bm{\alpha}}(\bm{x})=\text{Tr}[\rho_{H}(\bm{x})O(\mathbf{{\bm{\alpha}}})]$, using as training samples data of the form $\mathcal{T}^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{3}}=\{(\bm{x}_{\ell},y_{\ell})\}_{\ell=1}^{N}$ where $y_{\ell}\approx_{\epsilon_{3}}f^{\bm{\alpha}}(\bm{x}_{\ell})$ is an additive error $\epsilon_{3}$-approximation of the true expectation value $\text{Tr}[\rho_{H}(\bm{x}_{\ell})O({\bm{\alpha}})]$. Considering datasets with approximated values makes the learning problem closer to real world scenarios. In an idealized setting, the dataset would consist of pairs ${(\bm{x}_{\ell},y_{\ell})}$, where $y_{\ell}$ represents the exact expectation value. However, in real experiments, measurements are conducted by counting frequencies from a finite number of state copies, resulting in only approximate estimations of expectation values due to sampling errors. Formally, we assume a maximum (sampling) error $\epsilon_{3}$ on the training labels $y_{\ell}$ of our dataset, i.e. $\epsilon_{3}=\max_{\ell}|\text{Tr}[\rho_{H}(\bm{x}_{\ell})O({\bm{\alpha}})]-y_{\ell}|$. It is now possible to formally state our learning condition. Assuming the $\bm{x}_{\ell}$’s in the training data come from an unknown distribution $\mathcal{D}$, we require that the ML model learns the concept class $\mathcal{F}^{H,O}_{\mathrm{evolved}}$ in the following sense.
@@
A concept class $\mathcal{F}$ is efficiently learnable with respect to the distribution $\mathcal{D}$ if for every $\epsilon,\epsilon_{3}>0$ and $10<\delta<$ 1, there exists a poly$(1/\epsilon,1/\delta,1/\epsilon_{3},n)$-time algorithm $A$ such that for any $f^{\bm{\alpha}}$ in $\mathcal{F}$ and for any training set $\mathcal{T}^{\bm{\alpha}}_{\epsilon_{3}}$ of poly$(1/\epsilon,1/\delta,1/\epsilon_{3},n)$ size, $h(.)=A(\mathcal{T}_{\epsilon_{3}}^{\bm{\alpha}},.)$ satisfies with probability $1-\delta$:
@@
Notice that the learning algorithm possesses no prior knowledge regarding the observable $O({\bm{\alpha}})$, except for the fact that it is a $k$-local operator. Any additional information regarding $O({\bm{\alpha}})$ can exclusively be derived from the training samples within $\mathcal{T}^{\bm{\alpha}}_{\epsilon_{3}}$. In particular the vector ${\bm{\alpha}}$, which defines the specific concept in the concept class, is unknown to the learning algorithm. While the classical hardness of the learning problem relies on considering a specific ${\bm{\alpha}}$ for which the evaluation of $f^{\bm{\alpha}}(\bm{x})$ is known to be hard, the quantum algorithm will use a LASSO regression to infer a parameter $w\in[-1,1]^{m}$ close to the target ${\bm{\alpha}}$ so that condition (3) is satisfied.
@@
We emphasize here that our learning definition demands that the trained classical model can label new points. This stands in contrast to other settings, e.g. Hamiltonian learning problems, where the task would be identifying the vector ${\bm{\alpha}}$. In the prior case, a learning advantage is more easily established as explained in [GD23] where the hardness of evaluating versus identifying a concept was discussed. Later, we will explore this difference further and present an example of an identification problem closely related to the concept class of Eq. (1), which indeed can be solved by a classical algorithm.
@@
We can now state the first result of this paper, namely the existence of a concept class for the time evolution learning problem learnable by a quantum algorithm but for which no classical algorithm can meet the learning condition of Eq. (3).
Theorem 1 (Learning advantage for the time evolution problem).
@@
Under the assumption that there exists a distribution $\mathcal{D}^{\prime}$ such that $\mathrm{BQP}\not\subset\mathrm{HeurP}^{\mathcal{D}^{\prime}}/\mathrm{poly}$ [^1] We define the class $\mathrm{HeurP}^{\mathcal{D}}/\mathrm{poly}$ as the class of languages $\mathcal{L}$ that can be decided in $\mathrm{HeurP}/\mathrm{poly}$ under the distribution $\mathcal{D}$, i.e. $\mathrm{HeurP}^{\mathcal{D}}/\mathrm{poly}=\{\mathcal{L}\;\;|\;\;(\mathcal{L},\mathcal{D})\in\mathrm{HeurP}/\mathrm{poly}\}$ , then for any $\mathrm{BQP}$-complete language there exists an associated distribution $\mathcal{D}$ over the input values $\bm{x}\in\{0,1\}^{n}$, a Hamiltonian $H_{\mathrm{hard}}$ and a set of observables $\{O({\bm{\alpha}})\}_{\bm{\alpha}}$ such that no classical algorithm can solve the time evolution learning problem, formalized by the concept class $\mathcal{F}_{\mathrm{evolved}}^{H_{\mathrm{hard}},O}$, in the sense of Def. 1 given data $\mathcal{T}^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{3}}=\{(\bm{x}_{\ell},y_{\ell}\approx_{\epsilon_{3}}f^{\bm{\alpha}}(\bm{x}))\}_{\ell=1}^{N}$ sampled from $\mathcal{D}$. However, there exists a quantum algorithm which learns $\mathcal{F}_{\mathrm{evolved}}^{H_{\mathrm{hard}},O}$ under any distribution $\mathcal{D}$.

### v2

@@ Abstract
Quantum computers are believed to bring computational advantages in simulating quantum many body systems. However, recent works have shown that classical machine learning algorithms are able to predict numerous properties of quantum systems with classical data. Despite various examples of learning tasks with provable quantum advantages being proposed, they all involve cryptographic functions and do not represent any physical scenarios encountered in laboratory settings. In this paper we prove quantum advantages for the physically relevant task of learning quantum observables from classical (measured out) data. We consider two types of observables: first we prove a learning advantage for linear combinations of Pauli strings, then we extend the result for a broader case of unitarily parametrized observables. For each type of observable we delineate the boundaries that separate physically relevant tasks which classical computers can solve using data from quantum measurements, from those where a quantum computer is still necessary for data analysis. Differently from previous works, we base our classical hardness results on the weaker assumption that $\operatorname{\mathsf{BQP}}$ hard processes cannot be simulated by polynomial-size classical circuits and provide a non-trivial quantum learning algorithm. Our results shed light on the utility of quantum computers for machine learning problems in the domain of quantum many body physics, thereby suggesting new directions where quantum learning improvements may emerge.
I Introduction
@@
The very first proposed application of quantum computers can be traced back to Feynman’s idea of simulating quantum physics on a quantum device. Together with factoring [1], simulation of quantum many body systems stands as the clearest example of dramatic advantages of quantum computers [2]. Machine learning is another much newer area where quantum computers are believed to possibly bring advantages in certain learning problems, and in fact there are provable speed ups achieved by a quantum algorithm [3, 4, 5, 6, 7] for specific machine learning problems. Relating back to the original Feynman’s idea of simulating quantum physics, machine learning problems in quantum many body physics seem a natural scenario where learning advantages could arise. However, perhaps surprisingly, it was recently shown that access to data exemplifying what the hard-to-compute function does can drastically change the hardness of the computational task questioning the role of quantum computation in a machine learning scenarios [8, 9, 10]. If every quantum computation could be replicated classically provided access to data, such results would confine the practical application of quantum computers solely to the data acquisition stage. This is however not the case, as was shown already in the examples considered in [4, 3]. In such cases the unknown function was cryptographic in nature, and not related to genuine quantum simulation problems. Nonetheless, [11] introduced the first methods for establishing learning separations beyond cryptographic tasks. In particular, they demonstrated that learning problems with provable speed-ups can be constructed from any $\operatorname{\mathsf{BQP}}$-complete function, thereby enabling stronger links to physical scenarios. However, the work in [11] had two notable shortcomings. First, the classical non-learnability results were based on relatively strong assumptions about distributional problems, which, while plausible, are less well-understood than their decision problem counterparts. Second, the study did not introduce any significant quantum learning algorithms for general settings. Instead, quantum learnability was only shown in somewhat artificial scenarios where the concept class was small (polynomially bounded), allowing for the application of straightforward brute-force learning methods.
The results presented in this work effectively address both issues, enabling the establishment of clear boundaries between classical and quantum learning algorithms when dealing with data generated by quantum processes. Specifically, we consider learning problems where one is interested in predicting expectation values of an unknown observable from measurements on input quantum states which can either be ground states of local Hamiltonians or time evolved states. As motivation for the learning task, we have in mind experimentally plausible settings where learning an unknown observable may arise, such as in phase classification with an unknown order parameter or when dealing with real devices where the implemented measurement may be influenced by noise or external factors. In this scenarios, our result is a proof of learning advantages for different types of quantum observables. The main contributions of this work are as follows:
  • We prove an exponential quantum advantage in learning observables that are formed as a linear combination of local Pauli strings acting on time-evolved or ground states of local Hamiltonians. Importantly, by considering more general learning settings and leveraging results from classical learning theory, we base all of our learning advantage results on the more widely studied and generally accepted assumption that $\operatorname{\mathsf{BQP}}$-hard processes cannot be simulated by polynomial-size classical circuit (i.e., $\operatorname{\mathsf{BQP}}\not\subseteq\mathsf{P/poly}$), rather than on the less well-understood assumptions related to distributional problems.
  • Assuming $\operatorname{\mathsf{BQP}}\not\subseteq\mathsf{P/poly}$, we show how to construct learning problems that provide a quantum learning advantage in the general case of learning unitarily-parametrized observables, whenever an efficient procedure for learning unknown class of unitaries through query access exists. We then provide a concrete example by connecting to recent results on learning shallow unitaries.
We also examine the Hamiltonian learning problem where the identification of the target function is demonstrated to be classically easy. We summarize the learning settings and the achieved separations in Table 1, and provide practical examples of where these learning settings may arise in Section VI.
@@
Linear combination of Pauli strings Time-Evolution Problem Yes
@@
Table 1: Learning problems investigated in this paper.
II Related work
Although we are not the first to prove learning advantages for $\operatorname{\mathsf{BQP}}$-complete problems, our results present significant differences compared to previous works.
@@ II Related work
In [11] the authors first suggested that many learning speed-ups could arise by looking at $\operatorname{\mathsf{BQP}}$-complete problems. However, there are at least two crucial differences with our work. Firstly, the classical hardness result in [11] is derived by addressing heuristic complexity classes found in distributional problems. Specifically, it is assumed that there exists an input distribution such that a $\operatorname{\mathsf{BQP}}$-complete language $\mathcal{L}$ cannot be heuristically solved by a classical algorithm, i.e., $(\mathcal{L},\mathcal{D})\not\subseteq\mathsf{HeurP/poly}$. In contrast, in this work we employ a different proof technique and base all of our separation results on the assumption that $\operatorname{\mathsf{BQP}}\not\subseteq\mathsf{P/poly}$, thus focusing on the more widely investigated complexity classes of decision problems. Secondly, the learning problems considered in [11] have a concept class of polynomial size, and the quantum algorithm crucially leverages this to find the underlying concept in polynomial time by brute force. In this work we take a step further from their result by considering a physically relevant instantiation of a $\operatorname{\mathsf{BQP}}$-complete function. The associated concept classes are continuous and the quantum learning algorithm learns the unknown concept from data through a natural machine learning method.
@@
In recent works by [9, 10], significant progresses have been made with classical machine learning algorithms with provable guarantees for learning ground state properties from data, restricting the role of a quantum computer only to the data acquisition phase. In this study, we establish rigorous limitations on the capabilities of classical algorithms when learning quantum observables. A key distinction from their approach is that, in considering ground state problems, we investigate Hamiltonians with a polynomially decaying spectral gap, whereas the algorithms in [9, 10] have provable guarantees only for gapped Hamiltonians. The works in [12, 13] significantly extended the set of classical learnable quantum states to states which lies in the same Lindbladian phase. With respect to their works, our findings establish a definitive boundary on what can be learned through classical methods.
@@
In [14] the authors proposed a family of supervised learning tasks which present a learning speed-up with an exponentially sized concept class. Our construction, which was developed independently from [14], stems from a physically-motivated context, resulting in the utilization of different techniques for the quantum learning algorithm given the necessity to handle errors in the training samples. Furthermore, their complexity results are, once again, derived by focusing on distributional problems rather than decision problems.
Finally, in [15], the authors provide bounds on the overhead in training samples required by a classical algorithm compared to a quantum algorithm for the task of predicting a fixed observable. In contrast, our results focus primarily on the time efficiency of the learning algorithm. Furthermore, the results in [15] assume a fixed target distribution, whereas our setting requires learnability across all input distributions.
III Preliminaries
To make our work more self-contained and accessible, we include the necessary preliminaries from learning theory and complexity theory.
III.1 PAC learning
The definition of learnability in this work aligns directly with the widely adopted probably approximately correct (PAC) learning framework [16, 17]. In the case of supervised learning, a learning problem in the PAC framework is defined by a concept class $\mathcal{F}$ which, for each input size $n\in\mathbb{N}$, represent a set functions (or concepts) defined from some input space $\mathcal{X}_{n}$ (in this paper we assume $\mathcal{X}_{n}$ to be $\{0,1\}^{n}$ ) to some label set $\mathcal{Y}_{n}$ (in this paper we assume $\mathcal{Y}_{n}$ to be the interval $[-1,1]$). The learning algorithm receives as input information on the unknown target concept $f$ through training samples $T=\{(x_{\ell},f(x_{\ell}))\}_{\ell}$, where $\bm{x}\in\mathcal{X}_{n}$ is drawn according to a target distributions $\mathcal{D}_{n}$ over the inputs in $\mathcal{X}_{n}$. Finally, the goal of the learning algorithm is to output the description of a function (or hypothesis) $h$, which is again a function from $\mathcal{X}_{n}$ to $\mathcal{Y}_{n}$, that is in some sense “close” to the target concept $f$ which labels the data in $T$.
Definition 1 (Efficient PAC learnability).
A concept class $\mathcal{F}$ is efficiently PAC learnable if there exists a poly$(1/\epsilon,1/\delta,n)$-time algorithm A such that for all $\epsilon\geq 0$, all $0\leq\delta\leq 1$, and for any target concept $f$ in $\mathcal{F}$ and any target distribution $\mathcal{D}_{n}$ on $X_{n}$, if $A$ receives in input a training set $\mathcal{T}=\{(\bm{x}_{\ell},f(\bm{x}_{\ell}))\}_{\ell}$ of poly$(1/\epsilon,1/\delta,n)$ size, then with probability at least $1-\delta$ over the random samples in $\mathcal{T}$ and over the internal randomization of $A$, the learning algorithm $A$ outputs a specification of some hypothesis $h(.)=A(T,\epsilon,\delta,.)$ that satisfies
$\mathsf{Pr}_{\bm{x}\sim\mathcal{D}_{n}}\big{|}h(\bm{x})-f(\bm{x})\big{|}\leq\epsilon.$ (1)
In the above definition, learnability is required for any input distribution. While this might seems a strong requirement, it must be the case in order to ensure the applicability of the theory in real scenarios, where input distributions are often unknown or, in any case, not fixed.
III.2 Complexity theory
Our hardness result for the learning tasks discussed in this work is derived based on the complexity-theoretic assumption that $\operatorname{\mathsf{BQP}}\not\subseteq\mathsf{P/poly}$. We remark here that, in contrast to previous works [11], we do not consider the heuristic versions of the classes. Rather, we establish our learning separation by considering their simpler exact decision problem versions. In this section we provide the precise definition of this two complexity classes.
Definition 2 (BQP).
A language $\mathcal{L}$ is in BQP if and only if there exists a polynomial-time uniform family of quantum circuits $\{U_{n}:n\in\mathbb{N}\}$, such that
  1. 1.
    For all $n\in\mathbb{N}$, $U_{n}$ takes as input an $n$-qubit computational basis state, and outputs one bit obtained by measuring the first qubit in the computational basis.
  2. 2.
    For all $\bm{x}\in\mathcal{L}$, the probability that the output of $U_{|x|}$ applied on the input $\bm{x}$ is $1$ is greater or equal to $2/3$.
  3. 3.
    For all $\bm{x}\notin\mathcal{L}$, the probability that the output of $U_{|\bm{x}|}$ applied on the input $\bm{x}$ is 00 is greater or equal to $2/3$.
The class $\mathsf{P/poly}$, instead, captures the class of decision problems solvable by a polynomial time deterministic algorithm equipped with an “advice” bitstring. Importantly, the advice depends only on the input size, but it must be otherwise the same for every input $\bm{x}$ of a given size.
Definition 3 (Polynomial advice [18]).
A problem :$\mathcal{L}:\{0,1\}^{*}\rightarrow\{0,1\}$ is in $\mathsf{P/poly}$ if there exists a polynomial-time classical algorithm $\mathcal{A}$ with the following property: for every $n$ there exists an advice bitstring $\alpha_{n}\in\{0,1\}^{\mathrm{poly}(n)}$ such that for all $\bm{x}\in\{0,1\}^{n}$:
$\mathcal{A}(\bm{x},\alpha_{n})=\mathcal{L}(x).$ (2)
Equivalently, the class $\mathsf{P/poly}$ can be thought as the class of decision problem solvable by a non-uniform family of polynomial-size Boolean circuit. That is, for each input size, there exists an efficient circuit that correctly decides the inputs, but the circuits may differ completely for each input size $n$. Finally let us comment on the assumption $\operatorname{\mathsf{BQP}}\not\subseteq\mathsf{P/poly}$. Although the class $\mathsf{P/poly}$ is regarded as very powerful (for instance, it contains undecidable unary languages), it is generally believed that $\operatorname{\mathsf{BQP}}$ is not entirely contained within it. For example, if $\operatorname{\mathsf{BQP}}\subseteq\mathsf{P/poly}$, then factoring and the discrete logarithm problem [1] would also be solvable by non-uniform polynomial size classical circuits, which would compromise much of modern cryptography under standard security assumptions [19, 11]. Another argument against $\operatorname{\mathsf{BQP}}\subseteq\mathsf{P/poly}$ arises when considering the corresponding sampling problem. If quantum sampling could be done in $\mathsf{SampBPP/poly}$, then the polynomial hierarchy would collapse at the fourth level, as discussed in [20, 11, 21]. Finally, while Adleman’s trick allows classical randomness to be simulated with random strings, no analogous method exists for quantum algorithms. In particular, there is no known way to “extract the quantumness” [21] from a quantum algorithm, leaving the nature of polynomial advice for simulating quantum computers entirely unclear.
IV The learning problems
We now consider the first class of learning problems addressed in this paper, where the unknown observable is a linear combination of local Pauli strings. Within this framework, we define the following three abstract learning scenarios. In the first case, the input states come from an arbitrary distribution and are time-evolved under a fixed local Hamiltonian before being measured by the (partially) unknown target observable. In the second scenario, the inputs are known local Hamiltonians that time-evolve a fixed initial state before being measured. In the third case, we again consider different Hamiltonians as inputs, but here the learning problem involves learning measurements on their ground states, rather than on time-evolved states. We anticipate that these abstract scenarios can represent a variety of realistic settings where noise or external factors constrain control over the input Hamiltonians and the implemented measurements, we discuss more on this in Section VI.
@@ IV The learning problems
For all the three scenarios, we rigorously model the learning tasks using concept classes within the PAC learning framework introduced in Section III. We now analyze the first scenario, where the input states are time-evolved under a fixed Hamiltonian for a constant time, however, the input state preparation is not fully controlled. We model this by allowing the input states to be sampled randomly from an unknown distribution of quantum states, subsequently they undergo a time-evolution before being measured by an unknown observable. After collecting the corresponding measurement outcomes, the goal is to predict the expectation values of the unknown observable on new input states. Let $H$ be a Hamiltonian and fix a constant time $\tau$, we model the time-evolution learning problem by the following concept class
$\mathcal{F}_{\mathrm{evolved}}^{H,O}=\{f^{\bm{{\bm{\alpha}}}}(\bm{x})\in\mathbb{R}\;\;|\;\mathbf{{\bm{\alpha}}}\in[-1,1]^{m}\}$ (3)
with: $f^{\bm{\alpha}}(\bm{x}):\;\;\;\bm{x}\in\{0,1\}^{n}\rightarrow f^{\bm{\alpha}}(\bm{x})=\text{Tr}[\rho_{H}(\bm{x})O(\mathbf{{\bm{\alpha}}})]$
$O({\bm{\alpha}})=\sum_{i=1}^{m}{\bm{\alpha}}_{i}P_{i}.$
@@
In the above, $\bm{x}$ specifies the initial state |$\ket{\bm{x}}$, $\rho_{H}(\bm{x})$ the constant time evolved state $\rho_{H}(\bm{x})=U\ket{\bm{x}}\bra{\bm{x}}U^{\dagger}$ with $U=e^{iH\tau}$ and each $P_{i}$ is a $k$-local Pauli string where $m$ scales polynomially with $n$. Notice that although we consider binary inputs $\bm{x}\in\{0,1\}^{n}$, the output of each concept is real-valued. The goal of the ML learning algorithm is to learn a model $h(\bm{x})$ which approximates the unknown concept $f^{\bm{\alpha}}(\bm{x})=\text{Tr}[\rho_{H}(\bm{x})O(\mathbf{{\bm{\alpha}}})]$, using as training samples data of the form $\mathcal{T}^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{2}}=\{(\bm{x}_{\ell},y_{\ell})\}_{\ell=1}^{N}$ where $y_{\ell}\approx_{\epsilon_{2}}f^{\bm{\alpha}}(\bm{x}_{\ell})$ is an additive error $\epsilon_{2}$-approximation of the true expectation value $\text{Tr}[\rho_{H}(\bm{x}_{\ell})O({\bm{\alpha}})]$. Considering datasets with approximated values makes the learning problem closer to real world scenarios. In an idealized setting, the dataset would consist of pairs ${(\bm{x}_{\ell},y_{\ell})}$, where $y_{\ell}$ represents the exact expectation value. However, in real experiments, the estimations of these real-valued quantities are obtained by a finite number of state copies, resulting in only approximate estimations of expectation values due to sampling errors. Formally, we assume a maximum (sampling) error $\epsilon_{2}$ on the training labels $y_{\ell}$ of our dataset, i.e. $\epsilon_{2}=\max_{\ell}|\text{Tr}[\rho_{H}(\bm{x}_{\ell})O({\bm{\alpha}})]-y_{\ell}|$. It is now possible to formally state our learning condition. Assuming the $\bm{x}_{\ell}$’s in the training data come from an unknown distribution $\mathcal{D}$, we require that the ML model learns the concept class $\mathcal{F}^{H,O}_{\mathrm{evolved}}$ in the following sense  [^1] It is important to note that, in contrast to the learning condition in [11], here we require learnability under any distribution..
@@
A concept class $\mathcal{F}$ is efficiently learnable if there exists a poly$(1/\epsilon,1/\delta,1/\epsilon_{2},n)$-time algorithm $A$ such that for all $\epsilon,\epsilon_{2}>0$ and all $10<\delta<$ 1, and for any $f^{\bm{\alpha}}$ in $\mathcal{F}$ and any input target distribution $\mathcal{D}$, if $A$ receives in input a training set $\mathcal{T}^{\bm{\alpha}}_{\epsilon_{2}}$ of poly$(1/\epsilon,1/\delta,1/\epsilon_{2},n)$ size, $h(.)=A(\mathcal{T}_{\epsilon_{2}}^{\bm{\alpha}},\epsilon,\delta,.)$ satisfies with probability $1-\delta$:
@@
Notice that our learning condition directly follows from the definition of PAC learnability in Def. 1, with the only difference being that it is modified to account for errors in the training data, as discussed above.
@@
In our learning task, the learning algorithm only knows the observable through its functional form, mapping $\alpha$ to $O(\alpha)$. Any additional information regarding $O({\bm{\alpha}})$ can exclusively be derived from the training samples within $\mathcal{T}^{\bm{\alpha}}_{\epsilon_{2}}$. In particular the vector ${\bm{\alpha}}$, which defines the specific concept in the concept class, is unknown to the learning algorithm. While the classical hardness of the learning problem relies on considering a specific ${\bm{\alpha}}$ for which we prove the evaluation of $f^{\bm{\alpha}}(\bm{x})$ to be hard, the quantum algorithm will use a LASSO regression to infer a parameter $w\in[-1,1]^{m}$ close to the target ${\bm{\alpha}}$ so that condition (4) is satisfied.
@@
We emphasize here that our learning definition demands that the trained classical model can label new points. This stands in contrast to other settings, e.g. Hamiltonian learning problems, where the task would be identifying the vector ${\bm{\alpha}}$. In the prior case, a learning advantage is more easily established as explained in [11] where the hardness of evaluating versus identifying a concept was discussed. In Section V.2, we will explore this difference further and present an example of an identification problem closely related to the concept class of Eq. (3), which indeed can be solved by a classical algorithm.
@@
We can now state the first result of this paper, namely the existence of a concept class for the time-evolution learning problem learnable by a quantum algorithm but for which no classical algorithm can meet the learning condition of Eq. (4).
Theorem 1 (Learning advantage for the time-evolution problem).
@@
For any $\operatorname{\mathsf{BQP}}$-complete language there exists a Hamiltonian $H_{\mathrm{hard}}$ and a set of observables $\{O({\bm{\alpha}})\}_{\bm{\alpha}}$ such that no classical algorithm can efficiently solve the time-evolution learning problem, formalized by the concept class $\mathcal{F}_{\mathrm{evolved}}^{H_{\mathrm{hard}},O}$, in the sense of Def. 4, unless $\operatorname{\mathsf{BQP}}\subseteq\mathsf{P/poly}$. However, there exists a quantum algorithm which learns $\mathcal{F}_{\mathrm{evolved}}^{H_{\mathrm{hard}},O}$ under any input distribution $\mathcal{D}$.